OpenStudy (anonymous):

The perimeter of a basketball court is 90 meters and the length is 6 meters longer than twice the width. What are the length and the width?

OpenStudy (anonymous):

Use the length-width relation to restate the perimeter in terms of length and width.

OpenStudy (anonymous):

if length = 6 meters longer than twice the width, then L = 2W + 6 P = 2W + 2L 90 = 2W + 2(2W + 6) 90 = 6W + 12 78 = 6W W = 13 L=2(13)+6 L=32 W=13, L=32

OpenStudy (anonymous):

got it thanks